# Offline Imitation of Badminton Player Behavior via Experiential Contexts and Brownian Motion

### v1

@@ Abstract
In the dynamic and rapid tactic involvements of turn-based sports, badminton stands out as an intrinsic paradigm that requires alter-dependent decision-making of players. While the advancement of learning from offline expert data in sequential decision-making has been witnessed in various domains (e.g., video games), how to rally-wise imitate the behaviors of human players from offline badminton matches has remained underexplored. Replicating opponents’ behavior benefits players by allowing them to undergo strategic development with direction before matches. However, directly applying existing methods suffers from the inherent hierarchy of the match and the compounding effect due to the turn-based nature of players alternatively taking actions. In this paper, we propose RallyNet, a novel hierarchical offline imitation learning model for badminton player behaviors: (i) RallyNet captures players’ decision dependencies by modeling decision-making processes as a contextual Markov decision process. (ii) To facilitate the decision-making of an agent, RallyNet leverages the experience to generate context as the agent’s intent in the rally. (iii) To generate more realistic behavior, RallyNet leverages Geometric Brownian Motion (GBM) to model the interactions between players by introducing a valuable inductive bias for learning player behaviors. In this manner, RallyNet links player intents with interaction models with GBM, providing an understanding of real interactions for sports analytics. We extensively validate RallyNet with the largest available real-world badminton dataset consisting of men’s and women’s singles, demonstrating its ability to adeptly imitate player behaviors. The results illustrate RallyNet’s superiority, surpassing offline imitation learning methods and state-of-the-art turn-based approaches by at least 16% in the mean of rule-based agent normalization score. In addition, several practical use cases are discussed to showcase the applicability of RallyNet [^1] This study provides a comprehensive exposition of RallyNet, a component presented in the PhD symposium at the CIKM 2023 titled ”Enhancing Badminton Player Performance via a Closed-Loop AI Approach: Imitation, Simulation, Optimization, and Execution [1].” This study thoroughly details the implementation and experimental design for RallyNet, featuring specialized efficacy analyses customized for badminton scenarios.
Index Terms:
@@
Imitation learning, Inverse reinforcement learning, Badminton simulation, Sports analytics
I Introduction
Collecting historical data and simulating agents’ behaviors have been widely explored to study and replicate specific scenarios in various domains, e.g., autonomous driving [2], robots [3, 4] and gaming [5]. In sports analytics, one of the major goals is to understand and investigate the tactics of teams and individuals. Given the rapid development of information technology, massive amounts of fine-grained behavioral records in various sports domains have been collected to train models for analysis [6, 7, 8, 9]. On the other hand, these behavioral records also help models reproduce strategic behavior. The applicability to the replication of player strategic behavior could assist coaches in devising winning strategies. Moreover, it opens avenues for numerous applications, including tactics simulation [6] and sports broadcasting [10]. However, applying this learning to an online environment, particularly for badminton, a turn-based sport, poses unique challenges. Unlike basketball and football, where players dictate their own positions, in badminton, a player’s state (e.g., receiving position) is determined by the opponent. As a result, conventional imitation learning methods employed in basketball [11] and football [6] cannot be directly applied to turn-based sports since finding authentic opponents for the agent to interact with and enhance its strategies in a realistic game scenario is impractical.
@@ I Introduction
Therefore, we focus on leveraging offline behavioral records to replicate turn-based player behavior, employing offline imitation learning (IL) since it provides insights from past match data, enabling coaches and players to infer winning strategies. To effectively apply offline IL in badminton, we define three goals as key criteria to assess the model’s suitability for badminton scenarios.
@@
  • Rally Duration Realism: To accurately depict how actual players participate in matches, it is crucial for the learned agents to reproduce the genuine duration of rallies.
@@
Numerous past investigations have demonstrated success in offline IL, with Behavior Cloning (BC) standing out as a classical IL method for replicating a player’s actions. The effectiveness of offline IL has recently been illustrated, including hierarchical imitation learning (HIL) models [12, 13]. HIL models can excel in capturing the intrinsic hierarchy present in turn-based sports. For instance, in badminton, players often decide on a tactic before determining each action in a rally based on that strategy. Although hierarchical BC-based approaches enrich imitated expressivity in the long-horizontal task, none of them were designed for turn-based sports, which consist of multiple players taking actions alternatively to form a rally.
Therefore, there are two challenges to applying existing HIL methods directly for turn-based sports: 1) Leveraging experience. When players encounter situations that have appeared in their experience (e.g., from training and previous matches), they usually take corresponding actions for returning shots. Furthermore, in the same situation, there may be a wide variety of different corresponding behaviors in their experience. It is challenging to leverage experience to provide the helpful information of the action for the agent. 2) Alternative decision-making. In a turn-based sport, the state of each player is determined by the actions of not only themselves but also other players. Thus, the errors of an agent’s decision impact the decisions of other agents, resulting in more serious compounding errors. This challenge transcends the domain of multiple policies. It underscores how a player’s decision not only engages multiple policies but also directly determines the opponent’s state, shaping their decision dynamics. This stands in contrast to other multi-agent games such as board games, card games, and non-turn-based sports like basketball. In these games, a player’s actions do not directly dictate the opponent’s state since the actions decided by a player are not equivalent to the opponent’s state. This underscores the unique characteristic of turn-based sports.
@@
To address these challenges, we propose a hierarchical offline imitation learning model via experiential context and geometric Brownian motion (RallyNet) to capture long-term decision dependencies by modeling decision-making processes in turn-based sports as contextual Markov decision processes (CMDP) [14]. Based on the CMDP setting, the Experiential Context Selector (ECS) was designed by establishing the context space from experiences and selecting a context in the space as the agent’s intent to mimic the decision-making of the agents following their intents in the rally. This enables the agent’s behavior throughout the rally to not be influenced by partially incorrect decisions from the agent. If an incorrect decision occurs, adhering to the intent prevents it from spreading, enhancing decision reliability and providing the physical explanations of the context to understand the intent behind the agent’s decisions.
As the interactions between players are similar to those between particles, we introduce Latent Geometric Brownian Motion (LGBM) to capture the interactions between players. LGBM leverages the intrinsic capacity of Geometric Brownian Motion (GBM), characterized as a continuous-time stochastic process with Gaussian distributed increments [15], to model randomness and heavy-tailed distributions inherent in turn-based sports interactions. Players participating in a rally alternately complete GBM in latent space to enable the agent’s decision-making to jointly consider the opponent’s behavior. In this manner, RallyNet is able to generate more realistic behavior in alternative decision-making. Figure 1 illustrates how RallyNet is applied to a turn-based sport, and uses a badminton singles match as an example. An initial state of the rally including two players’ positions and score information is provided to recover the content of the rally.
We highlight our main contributions as follows:
  • We propose a novel HIL model via experiential context and geometric Brownian motion named RallyNet to learn player decision-making strategies in turn-based sports. To the best of our knowledge, this is the first work that imitates turn-based player behaviors.
  • RallyNet models turn-based decision-making processes as CMDP, and leverages experiences to generate agents’ intent, reducing the impact of partial incorrect decisions from the agent on overall behavior and enabling explanations to understand the intent behind the player’s behavior. In addition, we introduce the latent geometric Brownian motion to capture the interactions between players, making the generated behavior more realistic.
@@
II Preliminaries
II-A Badminton: A Typical Example of a Turn-Based Sport
@@
For simplicity, we consider a badminton game with two players (i.e., singles matches) as the demonstration example. The badminton environment has a real-world badminton court, which includes one player on each side, a shuttlecock, the net, and the boundary. As shown in Figure 1, the state of the player who hits the shuttlecock consists of the score information, the 2-dimensional position of the shuttlecock, the shot type to receive, the player’s position, the opponent’s position, and the moving vector of the opponent (e.g., move forward 1m). The action of the player consists of the landing position of the shuttlecock, the shot type (the shuttlecock type to hit), and the moving position to go to after returning the shuttlecock.
@@
II-B The Contextual Markov Decision Process
A Markov Decision Process (MDP) [16] is defined by a tuple $(S,A,T(y|x,a),R(x),\gamma)$, where $S$ is a set of states, $A$ is a set of possible actions agents can take, $T(y|x,a)$ is the transition probability, $R(x)$ is a reward function, and $\gamma\in[0,1)$ is the discount factor. At the $t$-th time step, the agent receives a state $s_{t}\in S$, then takes an action $a_{t}\in A$ according to a policy $\pi(a_{t}|s_{t})$. It is noted that the reward signal and the interaction with the environment are not available in the offline setting in this paper. The Contextual Markov Decision Process (CMDP) [14] leverages side information to extend the standard MDP with multiple contexts. A CMDP is defined by a tuple $(C,S,A,M)$, where $C$ is a set of contexts and $M$ is a function which maps a context $c\in C$ to MDP parameters $M(c)=\{R^{c},T^{c}\}$.
II-C Problem Formulation
For the badminton game with two players, we denote $P_{A}$ as the starting player and $P_{B}$ as the other player for each rally. Let =$\mathcal{R}=\{T_{r}\}^{|\mathcal{R}|}_{r=1}$ denote historical rallies of badminton matches, where the rally is composed of a sequence of state-action pairs. The $r$-th rally is denoted as $T_{r}=\{(s_{1}^{P_{A}},a_{1}^{P_{A}}),(s_{1}^{P_{B}},a_{1}^{P_{B}}),…,(s_{|T_{r}|}^{p},a_{|T_{r}|}^{p})\}$. At the $t$-th step, $s_{t}^{p}$ and $a_{t}^{p}$ represent the state and action of the player who takes action, respectively, where $p\in\{P_{A},P_{B}\}$ represents that the step is related to either $P_{A}$ or $P_{B}$. $p=P_{A}$ when $t=2i-1$ and $p=P_{B}$ when $t=2i$, where $i=1,2,\dots,|T_{r}|$. The action consists of the landing positions, shot type, and moving positions represented by $a_{t}^{p}=\langle l_{t}^{p},t_{t}^{p},m_{t}^{p}\rangle$. The imitation learning task for turn-based sports is to learn a policy that can recover players’ demonstrations from a set of historical rallies $\mathcal{R}$. Formally, for each demonstration rally in $\mathcal{R}$, given initial state $s_{1}^{a}$ of the $r$-th rally $T_{r}$, our goal is to recover the rally $T_{r}$.
Moreover, we define the experience of player $p$ with a current state $s_{t}^{p}$ as follows: For a rally that has experienced current states $s_{t}^{p}$, the experience is the action sequence of rallies $\{a_{t|p}^{p}\}$, where $t|p$ represents the step that is player $p$ taking action.
III Methodology
III-A Overview
Figure 2 illustrates the RallyNet framework, which consists of the following two components:
III-A1 Experiential Context Selector (Section III-C)
We substantiate the idea of leveraging experience via a two-step approach: (i) Experiential Context Selector (ECS) first extracts experiences by the experience extracting function; and (ii) ECS leverages experiences to construct a latent context space, which substantiates the concept of player intent by an encoder-decoder design, and provides a way to understand the intent behind the player’s behavior. Accordingly, ECS allows us to select a context as the agent’s selected intent of the rally, ensuring that its behavior during the rally persists unaffected by partially incorrect decisions.
III-A2 Latent Geometric Brownian Motion (Section III-D)
To better capture the interaction between players, we draw an analogy between particle motions and player motions in designing the Latent Geometric Brownian Motion module (LGBM), which frames players as particles to simulate the geometric Brownian motion in latent space. LGBM effectively brings its inductive bias into the player’s decisions and allows the model to consider the behavior of both players jointly instead of separately. Subsequently, the action projection layer (Section III-E) utilizes the latent position to predict both the landing and moving positions and the shot type for the next stroke. The algorithm is summarized in Algorithm 1.
Notably, we designed two processes in the training stage, namely the target process and the player process. The target process aims to learn to take actions that conform to the context by giving the correct actions as a target context for training. The player process utilizes experiences to learn how to select a context that approaches the output of target processes from the context space. Both processes share most of the modules, and only the player process is used during inferencing.
Algorithm 1 An Overview of RallyNet.
1:Historical rallies $\mathcal{R}$.
2:A policy which can recover players’ demonstration.
3:for each rally $r$ = 1 to $|\mathcal{R}|$ do
4:    for each step $t$ = 1 to $|T_{r}|$ do
5:         if $t$ == $1$ then
6:             $\hat{s}_{1}^{P_{A}}=s_{1}^{P_{A}}$
7:         else
8:             if $t\%2==0$ then
9:                 Decide next state of player B $\hat{s}_{t+1}^{P_{B}}$ by $\hat{a}_{t}^{P_{A}}$
10:             else
11:                 Decide next state of player A $\hat{s}_{t+1}^{P_{A}}$ by $\hat{a}_{t}^{P_{B}}$
12:             end if
13:         end if
14:         Get the state embedding of the player $x_{t}^{p}$ via Eq.1
15:         Collect experiences $\tau_{t}^{p}$ from $\mathcal{R}$ via Eq.2
16:         Select context $z_{ctx}$ from $\tau_{t}^{p}$ as intent via Eq.3-5
17:         $z_{t}^{\prime}=concat(x_{t}^{p},z_{ctx})$
18:         Get the latent position $z_{t}$ by $z_{t}^{\prime}$ and $z_{t-1}$ via Eq.10-11
19:         Project latent position $z_{t}$ into actions $\hat{a}_{t}^{p}$ via Eq.12-13
20:    end for
21:end for
III-B State Embedding Layer
To capture the long-term decision dependency, we concatenate the current state $s_{t}^{p}$ and historical states of the player who currently takes action $s_{1:t-1|p}^{p}$ and employ a transformer encoder to compute the embedding of the rally $r_{t}^{p}$.
To integrate the state with the corresponding player, a linear layer is adopted to compute the embedding of the player who hits the ball $e^{p}\in\mathbb{R}^{N_{p}}$ and concatenates with rally embedding $r_{t}^{p}\in\mathbb{R}^{d_{s}}$, where $d_{s}$ denotes the dimension of states and $N_{p}$ is the total number of players in the dataset. Formally, the output of the state embedding layer at the $t$-th step $x_{t}^{p}\in\mathbb{R}^{d_{s}\times N_{p}}$ is calculated as follows:
$\begin{split}r_{t}^{p}&=TRE(concat(s_{t}^{p},s_{1:t-1|p}^{p})),\\ x_{t}^{p}&=concat(W^{p}e^{p},r_{t}^{p}),\end{split}$ (1)
where $TRE$ denotes the transformer encoder and $W^{p}\in\mathbb{R}^{N_{p}}$ is a learnable matrix.
III-C Experiential Context Selector (ECS)

### v2

@@ Abstract
In the dynamic and rapid tactic involvements of turn-based sports, badminton stands out as an intrinsic paradigm that requires alter-dependent decision-making of players. While the advancement of learning from offline expert data in sequential decision-making has been witnessed in various domains (e.g., video games), how to imitate the behaviors of human players from rallies consisting of multiple players in offline badminton matches has remained underexplored. Replicating opponents’ behavior benefits players by allowing them to undergo strategic development with direction before matches. However, directly applying existing methods suffers from the inherent hierarchy of the match and the compounding effect due to the turn-based nature of players alternatively taking actions. In this paper, we propose RallyNet, a novel hierarchical offline imitation learning model for badminton player behaviors: (i) RallyNet captures players’ decision dependencies by modeling decision-making processes as a contextual Markov decision process. (ii) To facilitate the decision-making of an agent, RallyNet leverages the experience to generate context as the agent’s intent in the rally. (iii) To generate more realistic behavior, RallyNet leverages Geometric Brownian Motion (GBM) to model the interactions between players by introducing a valuable inductive bias for learning player behaviors. In this manner, RallyNet links player intents with interaction models with GBM, providing an understanding of real interactions for sports analytics. We extensively validate RallyNet with the largest available real-world badminton dataset consisting of men’s and women’s singles, demonstrating its ability to adeptly imitate player behaviors. The results illustrate RallyNet’s superiority, surpassing offline imitation learning methods and state-of-the-art turn-based approaches by at least 16% in the mean of rule-based agent normalization score. In addition, several practical use cases showcase the applicability of RallyNet.
Keywords:
@@
Imitation learning Inverse reinforcement learning Badminton simulation Sports analytics
1 Introduction
The exploration of simulating agents’ behaviors based on historical data has broad applicability across various domains. Whether it involves replicating specific scenarios for sports analytics [28], autonomous driving [3], or robots [25], these scenarios can be effectively framed as simulations of agents’ behaviors characterized by intricate interactions and decision-making processes. In sports analytics, one of the major goals is to understand and investigate the tactics of teams and individuals, which can be achieved by analyzing behavioral records to reproduce strategic behaviors. The applicability to the replication of player strategic behavior could assist in devising winning strategies [17] and sports broadcasting [11].
@@ 1 Introduction
However, unlike basketball and football, where players dictate their own positions, a player’s state (e.g., receiving position) is determined by the opponent in badminton, a typical turn-based sport. As a result, conventional imitation learning methods employed in basketball [30] and football [17] cannot be directly applied to turn-based sports since finding authentic opponents for the agent to interact with and enhance its strategies in a realistic game scenario is impractical. Therefore, we focus on leveraging offline behavioral records to replicate turn-based player behavior, employing offline imitation learning (IL) since it provides insights from past match data, enabling coaches and players to infer winning strategies. To effectively apply offline IL in badminton, three goals are defined as key criteria to assess the model’s suitability for badminton scenarios:
@@
  • Rally Duration Realism: To depict how actual players participate in matches, it is crucial for the agents to reproduce the duration of rallies.
@@
Recent investigations have showcased the success of offline imitation learning, including hierarchical imitation learning (HIL) models [31]. Although HIL approaches enrich imitated expressivity in the long-horizontal task, none of them were designed for turn-based sports, which consist of multiple players taking actions alternatively. Therefore, there are two challenges to applying existing HIL methods for turn-based sports: 1) Leveraging experience. When players encounter situations that have appeared in their experience (e.g., from training and previous matches), they usually take corresponding actions for returning shots. Furthermore, in the same situation, there may be a wide variety of different corresponding behaviors in their experience. It is challenging to leverage experience to provide the helpful information for the agent. 2) Alternative decision-making. In a turn-based sport, the state of each player is determined by the actions of not only themselves but also other players. Thus, The errors of one agent’s decisions affect others, leading to more severe compounded errors than in typical multi-agent imitation learning tasks [17].
@@
To address these challenges, we propose a hierarchical offline imitation learning model via experiential context and geometric Brownian motion (RallyNet) to capture long-term decision dependencies by modeling decision-making processes in turn-based sports as contextual Markov decision processes (CMDP) [13]. Based on the CMDP setting, the Experiential Context Selector (ECS) was designed by establishing the context space from experiences and selecting a context in the space as the agent’s intent to mimic the decision-making of the agents following their intents in the rally. This enables the agent’s behavior throughout the rally to not be influenced by partially incorrect decisions and provides the physical explanations of the context to understand the intent behind the player’s behavior. By viewing the interactions between players as those between particles, we introduce Latent Geometric Brownian Motion (LGBM) to capture the interactions between players. We make players participating in a rally alternately complete geometric Brownian motion [23] in latent space, which enables the agent’s decision-making to jointly consider the opponent’s behavior. This generates more realistic behavior in alternative decision-making. Fig. 1 illustrates how RallyNet is applied to a turn-based sport, and uses a badminton singles match as an example. An initial state of the rally including two players’ positions and score information is provided to recover the content of the rally. We highlight our main contributions as follows:
  • We propose a novel HIL model via experiential context and geometric Brownian motion named RallyNet to learn player decision-making strategies in turn-based sports.
  • RallyNet treats decision-making in badminton games as CMDPs, utilizing experiences to shape agents’ intent. Additionally, we introduce latent geometric Brownian motion to enhance realism by capturing player interactions.
@@
2 Related Work
2.1 Inverse Reinforcement Learning
Inverse Reinforcement Learning (IRL) [19] is an IL approach that attempts to learn the underlying rewards that an expert is optimizing from the demonstrations. With the learned reward function, IRL can guide the agent’s behavior toward that of the expert. However, in turn-based sports, each player’s actions determine the next state for other players. Treating opponents as part of the environment is equivalent to training in a constantly changing environment, which can make it challenging for IRL to learn optimal policies.
2.2 Offline Imitation Learning via Behavior Cloning
Behavior cloning (BC) provides a form of supervised learning for training policies by learning direct mapping from states to actions, which can be used in the offline setting [20]. Recent works such as DT [5], GCSL [10], and RvS [6] have shown that not only is supervised learning and bypassing the learning reward function able to attain better results for offline learning, but it is also easier to use and is more stable than offline inverse reinforcement learning (IRL) [19, 8]. Moreover, previous work [16, 31] has extended the BC to hierarchical BC by dividing a task into several sub-tasks, where the low-level behaviors are controlled by high-level decisions to capture long-term decision-making processes. However, these existing approaches focus on the complex behavior and the long-term task of the same target, but neglect the characteristics of the mixed sequences, which causes serious compounding errors.
3 Preliminaries
3.1 Badminton: A Typical Example of a Turn-Based Sport
@@
For simplicity, we consider a badminton game with two players (i.e., singles matches) as the demonstration example. As shown in Fig. 1, the state of the player who hits the shuttlecock consists of the score information, the 2-dimensional position of the shuttlecock, the shot type to receive, the player’s position, the opponent’s position, and the moving vector of the opponent (e.g., move forward 1m). The action of the player consists of the landing position of the shuttlecock, the shot type (the shuttlecock type to hit), and the moving position to go to after returning the shuttlecock.
3.2 The Contextual Markov Decision Process
A Markov Decision Process (MDP) [21] is defined by $(S,A,T(y|x,a),R(x),\gamma)$, where $S$ is a set of states, $A$ is a set of possible actions agents can take, $T(y|x,a)$ is the transition probability, $R(x)$ is a reward function, and $\gamma\in[0,1)$ is the discount factor. At the $t$-th time step, the agent receives a state $s_{t}\in S$, then takes an action $a_{t}\in A$ according to a policy $\pi(a_{t}|s_{t})$. It is noted that the reward signal and the interaction with the environment are not available in the offline setting in this paper. The Contextual Markov Decision Process (CMDP) [13] leverages side information to extend the standard MDP with multiple contexts. A CMDP is defined by a tuple $(C,S,A,M)$, where $C$ is a set of contexts and $M$ is a function which maps a context $c\in C$ to MDP parameters $M(c)=\{R^{c},T^{c}\}$.
3.3 Problem Formulation
For the badminton game with two players, we denote $P_{A}$ as the starting player and $P_{B}$ as the other player for each rally. Let =$\mathcal{R}=\{T_{r}\}^{|\mathcal{R}|}_{r=1}$ denote historical rallies of matches, where the rally is composed of a sequence of state-action pairs. The $r$-th rally is denoted as $T_{r}=\{(s_{1}^{P_{A}},a_{1}^{P_{A}}),(s_{1}^{P_{B}},a_{1}^{P_{B}}),…,(s_{|T_{r}|}^{p},a_{|T_{r}|}^{p})\}$. At the $t$-th step, $s_{t}^{p}$ and $a_{t}^{p}$ represent the state and action of the player who takes action, respectively, where $p\in\{P_{A},P_{B}\}$ represents that the step is related to either $P_{A}$ or $P_{B}$. $p=P_{A}$ when $t=2i-1$ and $p=P_{B}$ when $t=2i$, where $i=1,2,\dots,|T_{r}|$. The action consists of the landing positions, shot type, and moving positions represented by $a_{t}^{p}=\langle l_{t}^{p},t_{t}^{p},m_{t}^{p}\rangle$. The imitation learning task for turn-based sports is to learn a policy that can recover players’ demonstrations from a set of historical rallies $\mathcal{R}$. Formally, for each demonstration rally in $\mathcal{R}$, given initial state $s_{1}^{a}$ of the $r$-th rally $T_{r}$, our goal is to recover the rally $T_{r}$.
Moreover, we define the experience of player $p$ with a current state $s_{t}^{p}$ as follows: For a rally that has experienced current states $s_{t}^{p}$, the experience is the action sequence of rallies $\{a_{t|p}^{p}\}$, where $t|p$ represents the step that is player $p$ taking action. To speed up the extracting experience, we establish an experience extracting function $EXP$ to output the experience from the historical rallies corresponding to the current state.
We create a dictionary with keys as combinations of discrete values for player positions, opponent positions, shuttlecock positions (on a 10x10 grid), and shot types, and values as action sequences from rallies. By discretizing these continuous variables, we can efficiently retrieve matching rallies. When a new state is input, the function extracts the relevant positions and shot types, retrieves the corresponding action sequences of rallies from the dictionary, and then uses these action sequences to construct the context space.
@@ 3.3 Problem Formulation
4 Methodology
Fig. 2 illustrates the RallyNet framework consisting of two components: the Experiential Context Selector (ECS) and the Latent Geometric Brownian Motion (LGBM). The ECS component extracts experiences and constructs a latent context space to understand player intent, ensuring the agent’s behavior remains consistent during the rally. The LGBM component frames players as particles to simulate geometric Brownian motion, effectively capturing interactions between players and allowing joint consideration of their behavior. Notably, we designed two processes in the training stage, namely the target process and the player process. The target process aims to learn to take actions that conform to the context by giving the correct actions as a target context for training. The player process utilizes experiences to learn how to select a context that approaches the output of target processes from the context space. Both processes share most of the modules, and only the player process is used during inferencing. We proceed to present the proposed modules. In this paper, for a player $p$, we use $x_{t}^{p}$ to denote the state embedding for the $t$-th step of the rally. The state embedding is computed using a state embedding layer that processes the states experienced by the player through a Transformer encoder. The resulting embedding is then concatenated with the player embedding to obtain the final state embedding.
4.1 Experiential Context Selector (ECS)